OpenStudy (anonymous):

Clothes costs: 2 skirts $40 each 2 skirts $33 each Sales tax: 6 1/2% Shipping charge: $7.85 Total cost: ?

thomaster (thomaster):

(2*40)+(2*33)*1.065+7.85 So 80+66 = 146 you multiply that with 1.065 to add the sales tax 146*1.065=155.49 Now add the shipping charge 155.49+7.85= $163.34

thomaster (thomaster):

When the shipping charge is included in the tax, you add the 7.85 before you multiply with 1.065 So then it would be 146+7.85=153.85*1.065= $163.85

OpenStudy (texaschic101):

(2 * 40) + (2 * 33) = 80 + 66 = 146 6 1/2% = 6.5% = .065 146(.065) = 9.49 146 + 9.49 = 155.49 155.49 + 7.85 = 163.34 <-- total cost I do not think the shipping charge is included in the sales tax.